Session: Feb 2014 (session 1)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (C.Bortolotti at Fri Feb 28 15:00:59 2014) Minor failures:
Lost part of scan from 09:38:26 to 09:41:12 for antenna slewing - rfi in observed band

Noto (C. Nocita at Fri Feb 28 15:05:27 2014) Minor failures: HIGH
Tsys high from beginning of experiment to 9:50:22 due to low power output from local oscillator. Problem fixed not other known problems.

Westerbork (Geert Kuper at Fri Feb 28 15:23:15 2014) Success:

Yebes (Daniel Sastre Walter at Fri Feb 28 18:45:42 2014) Minor failures: WIND
Strong wind. We had to stop the experiment at 12:45 UTC.

Torun (K. Borkowski at Mon Mar 03 07:59:57 2014) Minor failures: FS
BBC07 unlocked throughout this experiment. Between 8:18 and 8:30 UTC telescope could have been off source due to hang-up of Field System.

Onsala (Jun Yang at Mon Mar 03 12:46:05 2014) Success:
Tsys increased significantly (2-4x) after 12:30 UT likely associated with the bad weather (rain).